# Contrastive Learning on Multimodal Analysis of Electronic Health Records

### v1

@@ 6 Discussion and Conclusion
In this paper, we propose the noisy multimodal log-linear production model for analyzing the multimodal EHR data and present the privacy-preserving algorithm CLAIME to estimate the multimodal feature embeddings with theoretical justification. Our algorithm is readily available for federated learning when multiple healthcare systems want to co-train their model but can not share patient-level data. It is also of great interest to consider a setting where different systems have overlapping but non-identical feature sets, in which case a block-wise matrix completion algorithm may be useful (Zhou et al.,, 2021).
Although our current analysis focuses on two modalities, it can be extended to more than two modalities. The rich EHR data contains many different modalities such as genetic data and image data. It will be interesting to extend our current framework to accommodate these data. Besides, it will also be interesting to incorporate the patient demographic in the generative model to consider the effects of variables like sex and age.
This paper focuses on a linear loss function, while non-linear loss functions such as mentioned in Remark 2.1 also deserve to be explored. Another interesting problem is the estimation of the patient embedding $\bm{c}_{i}$. The patient embedding is useful for many downstream tasks such as identifying “patients like me”.

### v2

@@ 6 Discussion and Conclusion
In this paper, we propose the noisy multimodal log-linear production model for analyzing multimodal EHR data and present the privacy-preserving algorithm CLAIME for estimating multimodal feature embeddings with theoretical justification. Our algorithm is readily applicable to federated learning when multiple healthcare systems wish to co-train their models without sharing patient-level data. It is also of great interest to consider settings where different systems have overlapping but non-identical feature sets, in which case block-wise matrix completion techniques may be useful (Zhou et al., , 2023).
While our main development focuses on the two-modality setting, CLAIME is conceptually extensible to more than two modalities, as illustrated in Supplementary S6. Theoretical justification, however, remains challenging. A central insight of our work is that, in the two-modality case, the loss function admits a closed-form global minimizer directly connected to the SVD of the PMI matrix. This reduction is crucial: it allows us to analyze the performance of the CLAIME representation through the top-$p$ singular space of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$. By contrast, the straightforward multimodal extension lacks such an explicit form, making the global minimizer analytically intractable.
This limitation reflects a broader challenge in the theory of contrastive learning, where rigorous analyses remain largely confined to the two-view case (Nakada et al., , 2023; Dufumier et al., , 2024). Moreover, our current extension assumes access only to pairwise co-occurrence matrices (Finlayson et al., , 2014; Beam et al., , 2020). If patient-level data were available, one could, in principle, construct higher-order co-occurrence tensors that capture joint dependencies across three or more modalities. Tensor SVD or related decompositions could then provide a natural way to generalize CLAIME embeddings with accompanying theoretical guarantees. Such a tensor-based framework could reveal interactions that are invisible to pairwise models, for example uncovering clinical pathways that involve coordinated use of lab tests, diagnoses, and treatments. Developing the statistical theory and scalable algorithms for such tensor-based generalizations remains an important direction for future work.
Beyond the linear loss considered here, exploring nonlinear losses such as those in Remark 2.4 is a natural next step. While a linear inner-product model captures only additive associations between modalities, nonlinear mappings (e.g., replacing the bilinear form $\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}$ with a neural network) could model richer dependencies. However, ensuring identifiability and statistical guarantees would require new theory beyond classical low-rank approximation, making the balance between flexibility, interpretability, and tractability a key open challenge. Another promising direction is the estimation of patient embeddings $\bm{c}_{i}$, which could enable clustering into clinically meaningful subgroups, modeling disease trajectories, and powering personalized applications such as “patients like me” recommendations. Such embeddings may be derived through joint factorization of patient–feature co-occurrence matrices or by incorporating temporal information from longitudinal data, with a theoretical framework for their statistical properties—consistency, stability, and generalization—being essential to ensure reliability in practice.